# Provably Faster Gradient Descent via Long Steps

### v1

@@ Abstract
This work establishes provably faster convergence rates for gradient descent via a computer-assisted analysis technique. Our theory allows nonconstant stepsize policies with frequent long steps potentially violating descent by analyzing the overall effect of many iterations at once rather than the typical one-iteration inductions used in most first-order method analyses. We show that long steps, which may increase the objective value in the short term, lead to provably faster convergence in the long term. A conjecture towards proving a faster $O(1/T\log T)$ rate for gradient descent is also motivated along with simple numerical validation.
@@ 1 Introduction
This work proposes a new analysis technique for gradient descent, establishing provably better convergence rates for smooth, convex optimization than the prior state-of-art textbook proofs. Our theory allows for nonconstant stepsize policies, periodically taking larger steps that may violate the monotone decrease in objective value typically needed by analysis. In fact, contrary to the common intuition, we show periodic long steps, which may increase the objective value in the short term, provably speed up convergence in the long term, with increasingly large gains as longer and longer steps are periodically included. This bears a similarity to accelerated momentum methods, which also depart from ensuring a monotone objective decrease at every iteration.
Establishing this requires new proof machinery capable of analyzing the overall effect of many iterations at once rather than the typical (naive) one-iteration inductions used in most first-order method analyses. Our proofs are based on the Performance Estimation Problem (PEP) ideas of [1, 2], which cast computing/bounding the worst-case problem instance of a given algorithm as a Semidefinite Program (SDP). We show that the existence of a feasible solution to a related SDP implies a convergence rate for gradient descent using a corresponding pattern of nonconstant stepsizes. This is in contrast to most existing PEP literature where computer-solves help guide the search for tighter convergence proofs [3, 4, 5, 6, 7, 8, 9, 10, 11, 12] and inform the development of new, provably faster algorithms [13, 14, 15, 16, 17]. Here the computer output itself constitutes the proof.
@@
given an initialization $x_{0}\in\mathbb{R}^{n}$. We assume throughout that a minimizer $x_{\star}$ of $f$ exists and its level sets are bounded $D=\sup\{\|x-x_{\star}\|_{2}\mid f(x)\leq f(x_{0})\}$ [^2] This assumption can likely be relaxed but is used to ease our development herein.. The classic convergence guarantee [18] for gradient descent is that with constant stepsizes $h=(1,1,1,\dots)$, every $T>0$ has
$f(x_{T})-f(x_{\star})\leq\frac{LD^{2}}{2(T+1)}\ .$ (1.2)
The proof of this result relies on the monotone decrease of the distance $\|x_{k}-x_{\star}\|_{2}$ and the objective $f(x_{k})$. Such monotone decreases are guaranteed if and only if $h_{k}\in(0,2)$. Taylor et al. [2] conjectured that this rate may be improvable by up to a factor of four by optimally tuning the constant stepsize used between one and two.
@@
for every even $T>0$. See our Theorem 3.2 characterizing many such alternating stepsize methods. The $+O(1/T^{2})$ term above is used to suppress two universal constants, namely we show there exist constants $\bar{s}$ and $C$ such that all even $T>2\bar{s}$ have bound $LD^{2}/(2.2\times T-C)$.
@@
Table 1: Improved convergence rates for Gradient Descent with stepsizes cycling through a “straightforward” pattern. Each convergence rate is proven by producing a certificate of feasibility for a related SDP, which is sufficient by our Theorems 2.1 and 3.1. Coefficients for $t\geq 7$ rates are slightly smaller than the ideal $\mathrm{avg}(h)$ due to rounding to produce an exact arithmetic certificate.
$x_{k+1}=x_{k}-\frac{h_{(k\mathtt{\ mod\ }t)}}{L}\nabla f(x_{k})\ .$ (1.3)
$f(x_{T})-f(x_{\star})\leq\frac{LD^{2}}{\mathrm{avg}(h)T}+O(1/T^{2})\ .$ (1.4)
See Section 2 for the formal introduction of this straightforwardness property. Hence the design of provably faster nonconstant gradient descent methods amounts to seeking straightforward stepsize patterns with large average stepsize values. Certifying a given pattern is straightforward can be done via semidefinite programming (see our Theorem 3.1). So the convergence rate analysis of such nonconstant stepsizes is a natural candidate for computer assistance. Table 1 shows straightforward stepsize patterns with increasingly fast convergence guarantees, each proven using a computer-generated, exact-arithmetic semidefinite programming solution certificate.
This work demonstrates that prior convergence theory and its focus on constant stepsizes and monotone objective decreases are insufficient to capture gradient descent’s potential. A first provably good path forward is offered.
@@
Future works identifying longer straightforward patterns and other tractable families of nonconstant, periodically long stepsize policies will surely be able to improve on this work’s particular rates.
@@
The analysis of such nonconstant, long stepsize gradient descent methods has eluded the literature, with one exception. In 1953, Young [19] showed optimal, accelerated convergence is possible for gradient descent when minimizing a smooth, strongly convex quadratic function by using a careful nonconstant selection of $h_{i}$. Namely, Young set $h_{0}\dots h_{T-1}$ as one over the roots of the $T$-degree Chebyshev polynomial [^3] A nice summary of this is given by the recent blog post [20].. No progress has been made beyond quadratics in accelerating gradient descent via long steps. This is partly due to Young’s arguments relying fundamentally on framing gradient descent as repeated matrix multiplication, lacking a clear path to generalize. The primary innovation in this work can then be viewed as identifying a tractable analysis technique capable of advancing in this seventy-year-old open direction.
@@
The search for long, straightforward stepsize patterns $h$ is hard; the set of all straightforward patterns is nonconvex, making local searches often unfruitful. Our patterns of length $t=2^{m}-1$ in Table 1 were created by repeating the pattern for $t=2^{m-1}-1$ twice with a new long step added in between and (by hand) shrinking the long steps in the length $12^{m-1}-$ 1 subpatterns. This doubling procedure consistently increased $\mathrm{avg}(h)$ by $\approx 0.6$. We conjecture the following.
@@
For any $t$, there exists a straightforward $h\in\mathbb{R}^{t}$ with $\mathrm{avg}(h)=\Omega(\log(t))$.
If true, this would likely yield convergence rates on the order of $O(1/T\log(T))$, strictly improving on the classic $O(1/T)$ guarantee. If such long patterns exist, one natural question is how close to the optimal $O(1/T^{2})$ rate attained by momentum methods can be achieved by gradient descent with long steps. The numerics of [12, Figure 2] suggested a $O(1/T^{1.178})$ rate may be possible. The theory of Lee and Wright [21] showed asymptotic $o(1/T)$ convergence for constant stepsize gradient descent, which also motivates the possibility for improved convergence rates like are presented here.
For strongly convex optimization (or, more generally, any problem satisfying a Hölder growth bound), the classic convergence rates for constant stepsize gradient descent are known to improve. We show the same improvements occur for any straightforward stepsize pattern (see Theorem 2.2) with an additional gain of $\mathrm{avg}(h)$. We validate that the convergence speed-ups of Table 1 actually occur on randomly generated least squares problems in Figure 1, seeing gains proportional to $\mathrm{avg}(h)$. These plots also showcase that descent is not ensured within the execution of a straightforward pattern as $t=7$ rapidly oscillates within each pattern while converging overall.
Figure 1: Least squares problems minimizing $\|Ax-b\|^{2}_{2}$ (left) and $\|Ax-b\|^{2}_{2}+\|x\|^{2}_{2}$ (right) with i.i.d. normal entries in $A\in\mathbb{R}^{n\times n}$ and $b\in\mathbb{R}^{n}$ for $n=4000$. Gradient Descent (1.3)’s objective gap is plotted over $T=2000$ iterations with $h=(1)$ and with each pattern from Table 1. Note this second objective is substantially more strongly convex, so its faster linear convergence is expected.
@@ 1.1 Sketch of Proof Technique - Reducing proving eventual descent to an SDP
For constant stepsizes equal to one, this amounts to $f(x_{0}-\nabla f(x_{0})/L)\leq f(x_{0})-\delta^{2}/LD^{2}$, a classic descent result that holds for all $L$-smooth convex $f$. To prove a descent lemma like (1.5), we consider taking a direct combination of several known inequalities. We are given the equalities $x_{k+1}=x_{k}-(h_{k}/L)\nabla f(x_{k})$ and $\nabla f(x_{\star})=0$ and as inequalities an initial distance bound
@@
and for any $x_{i}$ and $x_{j}$ with $i,j\in\{\star,0,1,2,\dots t\}$, convexity and smoothness imply[22, (2.1.10)]
@@
$+\sum_{i,j\in\{\star,0,\dots t\}}\lambda_{i,j}(\delta)$ $\left(f(x_{j})+\nabla f(x_{j})^{T}(x_{i}-x_{j})+\frac{1}{2L}\|\nabla f(x_{i})-\nabla f(x_{j})\|^{2}_{2}-f(x_{i})\right)\leq 0\ .$ (1.6)
Our proof then proceeds by showing carefully selected values of $w(\delta),v(\delta),\lambda(\delta)$ reduce this inequality to guaranteeing (1.5). We find to prove our rates in Table 1 it suffices to set $v(\delta)=\frac{\sum_{i=0}^{t-1}h_{i}}{D^{2}}\delta^{2}$, $w(\delta)=1-\frac{2\sum_{i=0}^{t-1}h_{i}}{LD^{2}}\delta$, and use a linear function $\lambda(\delta)=\lambda+\delta\gamma$. For this choice of $v(\delta)$ and $w(\delta)$, our Theorem 3.1 shows (1.6) implies (1.5) if $(\lambda,\gamma)$ is feasible to a certain SDP. Hence proving a $t$ iteration descent lemma of the form (1.5) can be done by semidefinite programming. Our Theorem 2.1 then completes the argument by showing such a periodic descent guarantee implies the rate (1.4).
@@ 2 Straightforward Stepsize Patterns
The key insight of Drori and Teboulle [1] is that this infinite-dimensional problem can be reformulated as a finite-dimensional semidefinite program. We formalize and utilize this insight in Section 3.
Generally, the worst-case functions $f$ attaining (2.1) can be nontrivial. To find a tractable family of stepsizes for analysis, we focus on ones where this worst-case behavior is no worse than a simple one-dimensional setting: Consider the one-dimensional (nonsmooth) convex function linearly decreasing from $x_{0}$ to $x_{\star}$ and constant thereafter. That is, given $L,D,\delta$, consider the problem instance with $f(x)=\max\{\delta x/D,0\}$, $x_{0}=D$, and $x_{\star}=0$. Provided $\delta$ is small enough (i.e., $\delta\leq LD^{2}/\sum h_{i}$), the gradient descent iteration $x_{k+1}=x_{k}-\frac{h_{k}}{L}f^{\prime}(x_{k})$ has
@@
We say that a stepsize pattern $h$ is straightforward if its worst-case behavior (over all smooth functions, as defined in (2.1)) is no worse than this one-dimensional piecewise linear setting for $\delta$ small enough. Formally, we say that a stepsize pattern $h$ is straightforward if for some $\Delta\in(0,1/2]$,
@@
for any $L,D>0$. Our analysis of gradient descent with periodic long steps proceeds by:
@@
The second step is simpler, so we address it first. Additionally, we show the same factor of rate improvement $\mathrm{avg}(h)$ carries over to more structured domains like strongly convex optimization.

### v5

@@ Abstract
This work establishes new convergence guarantees for gradient descent in smooth convex optimization via a computer-assisted analysis technique. Our theory allows nonconstant stepsize policies with frequent long steps potentially violating descent by analyzing the overall effect of many iterations at once rather than the typical one-iteration inductions used in most first-order method analyses. We show that long steps, which may increase the objective value in the short term, lead to provably faster convergence in the long term. A conjecture towards proving a faster $O(1/T\log T)$ rate for gradient descent is also motivated along with simple numerical validation.
@@ 1 Introduction
This work proposes a new analysis technique for gradient descent, establishing a path toward provably better convergence guarantees for smooth, convex optimization than is possible with existing approaches based on constant stepsizes. Instead, our theory allows for nonconstant stepsize policies, periodically taking larger steps that may violate the monotone decrease in objective value typically needed by analysis. In fact, contrary to the common intuition, we show periodic long steps, which may increase the objective value in the short term, provably speed up convergence in the long term, with increasingly large gains as longer and longer steps are periodically included. This bears a similarity to accelerated momentum methods, which also depart from ensuring a monotone objective decrease at every iteration.
Establishing this requires a proof technique capable of analyzing the overall effect of many iterations at once rather than the typical (naive) one-iteration inductions used in most first-order method analyses. Our proofs are based on the Performance Estimation Problem (PEP) ideas of [1, 2, 3], which cast computing/bounding the worst-case problem instance of a given algorithm as a Semidefinite Program (SDP). We show that the existence of a feasible solution to a related SDP proves a descent guarantee after applying a corresponding pattern of nonconstant stepsizes, from which faster convergence guarantees follow. Our technique is very similar to that first proposed by Altschuler’s Master’s thesis [4, Chapter 8] which established repeating stepsize patterns of length two or three with faster contractions towards the minimizer for smooth, strongly convex minimization. Most existing PEP literature uses computer-solves to guide the search for tighter convergence proofs [5, 3, 6, 7, 8, 9, 10, 11, 12, 13] and inform the development of new, provably faster algorithms [14, 15, 16, 17, 18]. Here computer outputs are directly used to constitute the proof, but are large (up to tens of thousands of rational numbers) and so may provide less guidance or intuition.
@@
given an initialization $x_{0}\in\mathbb{R}^{n}$. We assume throughout that a minimizer $x_{\star}$ of $f$ exists and its level sets are bounded $D=\sup\{\|x-x_{\star}\|_{2}\mid f(x)\leq f(x_{0})\}$ [^2] This assumption can likely be relaxed but eases our development herein.. The classic convergence guarantee [19] for gradient descent is that with constant stepsizes $h=(1,1,1,\dots)$, every $T>0$ has $f(x_{T})-f(x_{\star})\leq LD^{2}/2(T+1)$. Using performance estimation techniques, a tight refinement of this bound was given by [1, Theorem 3.1] (and elementary proof, avoiding the use of PEP, was given by [20]): every $T>0$ has
$f(x_{T})-f(x_{\star})\leq\frac{LD^{2}}{4T+2}\ .$ (1.2)
By tight, we mean a matching problem instance exists that attains the above inequality. Using stepsizes between one and two, this rate can be improved by another factor of two (see [2, 20]). Stepsizes beyond length two have very little prior theory as one can no longer guarantee a decrease in objective value at each iteration.
@@
for every even $T>0$. See our Theorem 3.2 characterizing many such alternating stepsize methods. The $+O(1/T^{2})$ terms throughout this work are only used to suppress two universal constants, namely above, we show there exist constants $\bar{s}$ and $C$ such that all even $T>2\bar{s}$ have bound $LD^{2}/(2.2\times T-C)$.
$x_{k+1}=x_{k}-\frac{h_{(k\mathtt{\ mod\ }t)}}{L}\nabla f(x_{k})\ .$ (1.3)
$f(x_{T})-f(x_{\star})\leq\frac{LD^{2}}{\mathrm{avg}(h)T}+O(1/T^{2})$ (1.4)
where $\mathrm{avg}(h)=\frac{1}{t}\sum_{i=0}^{t-1}h_{i}$. See Section 2 for the formal introduction of this straightforwardness property. Hence the design of provably faster nonconstant stepsize gradient descent methods amounts to seeking straightforward stepsize patterns with large average stepsize values. Certifying a given pattern is straightforward can be done via semidefinite programming (see our Theorem 3.1). So the convergence analysis of such nonconstant stepsizes methods is a natural candidate for computer assistance.
@@
Table 1: Improved convergence guarantees for Gradient Descent with stepsizes cycling through a “straightforward” pattern. Each convergence bound is proven by producing a certificate of feasibility for a related SDP, which is sufficient by our Theorems 2.1 and 3.1. Coefficients for $t\geq 7$ guarantees are slightly smaller than the ideal $\mathrm{avg}(h)$ due to rounding to produce an exact arithmetic certificate.
@@
Table 1 shows straightforward stepsize patterns with increasingly fast convergence guarantees, each proven using a computer-generated, exact-arithmetic semidefinite programming solution certificate. Future works identifying longer straightforward patterns and other tractable families of nonconstant, periodically long stepsize policies will surely be able to improve on this work’s particular guarantees.
@@
The analysis of such nonconstant, long stepsize gradient descent methods has eluded the literature, with only a few exceptions. In 1953, Young [21] showed optimal, accelerated convergence is possible for gradient descent when minimizing a smooth, strongly convex quadratic function by using a careful nonconstant selection of $h_{i}$. Namely, Young set $h_{0}\dots h_{T-1}$ as one over the roots of the $T$-degree Chebyshev polynomial [^3] A nice summary of this is given by the recent blog post [22].. Few works have shown faster convergence from long steps beyond quadratics. Two notable works have done so for smooth, strongly convex minimization:
Oymak [23] showed substantial speed-ups for strongly convex functions with special bimodal structured Hessians. Closely in line with this work’s reasoning, several recent Master’s and doctoral theses have addressed the optimal design and analysis of one, two, or three steps of gradient descent. The theses of Daccache [24] and Eloi [25] provide exhaustive characterizations of gradient descent’s objective gap after two or three steps for smooth convex optimization. Their results do not provide a mechanism to be applied inductively, so no convergence rates follow from repeatedly applying the studied patterns. For strongly convex optimization, Altschuler [4, Chapter 8] used PEP techniques to derive the optimal patterns of length $t=1,2,3$ for contracting either the distance to optimal or the gradient’s norm. Inductively applying their improved contractions, one arrives at the same moral takeaway advanced by this work: SDP-based analysis can prove faster convergence follows from periodically taking longer steps, potentially violating descent. Importantly, their length two and three patterns are optimal, meaning their patterns give the best possible contraction factor. Note their patterns differ from those presented here, only being optimal for strongly convex problems. The primary contribution in this work is identifying a tractable analysis technique for general smooth, convex optimization and showing increasing performance gains for increasingly large $t>3$, continuing in this seventy-year-old direction.
@@
The search for long, straightforward stepsize patterns $h$ is hard; the set of all straightforward patterns is nonconvex, making local searches often unfruitful. Our patterns of length $t=2^{m}-1$ in Table 1 were created by repeating the pattern for $t=2^{m-1}-1$ twice with a new long step added in between and (by hand) shrinking the long steps in the length $12^{m-1}-$ 1 subpatterns. This recursive pattern has strong similarities to the cyclic and fractal Chebyshev patterns for quadratic minimization considered by [26, 27, 28], although we make no provable connections. This doubling procedure consistently increased $\mathrm{avg}(h)$ by $\approx 0.6$. We conjecture the following.
@@
For any $t\in\mathbb{N}$, there exists a straightforward stepsize pattern $h\in\mathbb{R}^{t}$ with $\mathrm{avg}(h)=\Omega(\log(t))$.
If true, this would likely yield convergence rates on the order of $O(1/(T\log(T)))$, strictly improving on the classic $O(1/T)$ guarantee. If such long patterns exist, one natural question is how close to the optimal $O(1/T^{2})$ rate attained by momentum methods can be achieved by gradient descent with long steps. The numerics of [13, Figure 2] suggested a $O(1/T^{1.178})$ rate may be possible. The theory of Lee and Wright [29] showed asymptotic $o(1/T)$ convergence for constant stepsize gradient descent (although no improved guarantees in finite-time were given), which also motivates the possibility for improved finite-time convergence guarantees.
Figure 1: Least squares problems minimizing $\|Ax-b\|^{2}_{2}$ (left) and $\|Ax-b\|^{2}_{2}+\|x\|^{2}_{2}$ (right) with i.i.d. normal entries in $A\in\mathbb{R}^{n\times n}$ and $b\in\mathbb{R}^{n}$ for $n=4000$. Objective gaps are plotted over $T=2000$ iterations with $h=(1)$ and with each pattern from Table 1. Note this second objective is substantially more strongly convex, so its faster linear convergence is expected. For comparison, Nesterov’s accelerated method (without modification to utilize strong convexity) is shown.
For strongly convex optimization (or, more generally, any problem satisfying a Hölder growth bound), the classic convergence rates for constant stepsize gradient descent are known to improve. We show the same improvements occur for any straightforward stepsize pattern (see Theorem 2.2) with an additional gain of $\mathrm{avg}(h)$. We validate that the convergence speed-ups of Table 1 actually occur on randomly generated least squares problems in Figure 1, seeing gains proportional to $\mathrm{avg}(h)$. These plots also showcase that descent is not ensured within the execution of a straightforward pattern as $t=7$ rapidly oscillates within each pattern while converging overall. Note, gradient descent is not an optimal first-order method as accelerated methods can attain a faster $O(1/T^{2})$ rate. A basic implementation of Nesterov’s accelerated method is included in our numerics to showcase the degree of progress in closing this gap able to be accomplished via our longer steps.
Cyclic, periodically long stepsizes have been used in neural network training schedules [30, 31, 32]. Our results only apply to convex, deterministic problems. New techniques handling nonconvexity and stochasticity would need to be developed to describe the effect of long steps in such settings of machine learning.
@@ 1.1 Sketch of Proof Technique - Reducing proving eventual descent to an SDP
For constant stepsizes equal to one, this amounts to $f(x_{0}-\nabla f(x_{0})/L)\leq f(x_{0})-\delta^{2}/LD^{2}$, a classic descent result that holds for all $L$-smooth convex $f$.
To prove a descent lemma like (1.5), one can take a direct combination of several known inequalities. This is a well-known approach, equivalent to providing dual solutions to the dual of the associated performance estimation problem. We are given the equalities $x_{k+1}=x_{k}-(h_{k}/L)\nabla f(x_{k})$ and $\nabla f(x_{\star})=0$ and as inequalities an initial distance bound
@@
and for any $x_{i}$ and $x_{j}$ with $i,j\in\{\star,0,1,2,\dots t\}$, convexity and smoothness imply[33, (2.1.10)]
@@
$+\sum_{i,j\in\{\star,0,\dots t\}}\lambda_{i,j}(\delta)$ $\left(f(x_{j})+\nabla f(x_{j})^{T}(x_{i}-x_{j})+\frac{1}{2L}\|\nabla f(x_{i})-\nabla f(x_{j})\|^{2}_{2}-f(x_{i})\right)\leq 0\ .$
Our proof then proceeds by showing carefully selected functions $w(\delta),v(\delta),\lambda(\delta)$ reduce this inequality to guaranteeing (1.5). We find to prove our guarantees in Table 1, it suffices to set $v(\delta)=\frac{\sum_{i=0}^{t-1}h_{i}}{LD^{4}}\delta^{2}$, $w(\delta)=1-\frac{2\sum_{i=0}^{t-1}h_{i}}{LD^{2}}\delta$, and use a linear function $\lambda(\delta)=\lambda+\delta\gamma$. For this choice of $v(\delta)$ and $w(\delta)$, our Theorem 3.1 shows (1.6) implies (1.5) if $(\lambda,\gamma)$ is feasible to a certain SDP. Hence, proving a $t$ iteration descent lemma of the form (1.5) can be done by semidefinite programming. Our Theorem 2.1 then completes the argument by showing such a periodic descent guarantee implies the guarantee (1.4).
@@ 2 Straightforward Stepsize Patterns
Note this problem is infinite-dimensional as it optimizes over functions $f\colon\mathbb{R}^{n}\rightarrow\mathbb{R}$ and vectors $x_{0},x_{\star}\in\mathbb{R}^{n}$ for some given problem dimension $n$. The key insight of Drori and Teboulle [1] is that such infinite-dimensional problems can be relaxed to a finite-dimensional semidefinite program. Subsequently, Taylor et al. [2] showed this relaxation can be made an exact reformulation. We formalize and utilize these insights in Section 3. We note one particular nice property here: $p_{L,D}$ is concave.
Lemma 2.1.
For any $L,D>0$, $p_{L,D}$ is concave.
Proof.
Consider any $\delta^{(1)},\delta^{(2)}\in\mathbb{R}$ and $\lambda\in(0,1)$. Denote feasible solutions to the problem defining $p_{L,D}(\delta^{(i)})$ for each $i=1,2$ by $f^{(i)},x^{(i)}_{0},x^{(i)}_{\star}$. Note if no feasible solution exists for either of these problems, one vacuously has concavity as
$\lambda p_{L,D}(\delta^{(1)})+(1-\lambda)p_{L,D}(\delta^{(2)})=-\infty\leq p_{L,D}(\lambda\delta^{(1)}+(1-\lambda)\delta^{(2)})\ .$
Note without loss of generality, $x^{(i)}_{\star}=0$ and $f^{(i)}(x^{(i)}_{\star})=0$. Then one can easily check $\lambda f^{(1)}+(1-\lambda)f^{(2)},\lambda x_{0}^{(1)}+(1-\lambda)x_{0}^{(2)},\lambda x_{\star}^{(1)}+(1-\lambda)x_{\star}^{(2)}$ is feasible to the problem defining $p_{L,D}(\lambda\delta^{(1)}+(1-\lambda)\delta^{(2)})$. Consequently, taking the supremum over choices of $f^{(i)},x_{0}^{(i)}$ gives concavity as
$\lambda p_{L,D}(\delta^{(1)})+(1-\lambda)p_{L,D}(\delta^{(2)})\leq p_{L,D}(\lambda\delta^{(1)}+(1-\lambda)\delta^{(2)})\ .$
Generally, the worst-case functions $f$ attaining (2.1) can be nontrivial. To find a tractable family of stepsizes for analysis, we focus on ones where this worst-case behavior is no worse than a simple one-dimensional setting: Consider the one-dimensional (nonsmooth) convex function linearly decreasing from $x_{0}$ to $x_{\star}$ and constant thereafter. That is, given $L,D,\delta$, consider the problem instance $f(x)=\max\{\delta x/D,0\}$, $x_{0}=D$, and $x_{\star}=0$. Provided $\delta$ is small enough (i.e., $\delta\leq LD^{2}/\sum_{i=0}^{t-1}h_{i}$), the gradient descent iteration $x_{k+1}=x_{k}-\frac{h_{k}}{L}\nabla f(x_{k})$ has
@@
We say that a stepsize pattern $h$ is straightforward if its worst-case behavior (over all smooth functions, as defined in (2.1)) is no worse than this one-dimensional piecewise linear setting for $\delta$ small enough. Formally, we say that a stepsize pattern $h$ of length $t$ is straightforward if for some $\Delta\in(0,1/2]$,
@@
for any $L,D>0$. Our analysis of gradient descent with long steps then proceeds by:
@@
The second step is simpler, so we address it first. Additionally, we show the same factor of guarantee improvement $\mathrm{avg}(h)$ carries over to more structured domains like strongly convex optimization.